# When Audits and Recounts Distract from Election Integrity: The 2020 U.S. Presidential Election in Georgia

### v1

@@
Abstract
The U.S. state of Georgia was central to efforts to overturn the results of the 2020 Presidential election, including a phone call from then-president Donald Trump to Georgia Secretary of State Brad Raffensperger asking Raffensperger to ‘find’ 11,780 votes. Raffensperger has maintained that a ‘100% full-count risk-limiting audit’ and a machine recount agreed with the initial machine-count results, which proved that the reported election results were accurate and that ‘no votes were flipped.’ While there is no indication of widespread fraud, there is reason to distrust the election outcome: the two machine counts and the manual ‘audit’ tallies disagree substantially, even about the number of ballots cast. Some ballots in Fulton County, Georgia, were included in the original count at least twice; some were included in the machine recount at least thrice. Audit handcount results for some tally batches were omitted from the reported audit totals: reported audit results do not include every vote the auditors counted. In short, the two machine counts and the audit were not probative of who won because of poor processes and controls: a lack of secure physical chain of custody, ballot accounting, pollbook reconciliation, and accounting for other election materials such as memory cards. Moreover, most voters voted with demonstrably untrustworthy ballot-marking devices, so even a perfect handcount or audit would not necessarily reveal who really won. True risk-limiting audits (RLAs) and rigorous recounts can limit the risk that an incorrect electoral outcome will be certified rather than being corrected. But no procedure can limit that risk without a trustworthy record of the vote. And even a properly conducted RLA of some contests in an election does not show that any other contests in that election were decided correctly. The 2020 U.S. Presidential election in Georgia illustrates unrecoverable errors that can render recounts and audits ‘security theater’ that distract from the more serious problems rather than justifying trust.
@@ Abstract
Risk-limiting audit, recount, evidence-based elections, U.S. presidential election
1 Introduction: The 2020 U.S. Presidential Election in the State of Georgia
Georgia was one of the ‘swing states’ that determined the outcome of the 2020 U.S. presidential election: its 16 electoral college votes went to Joe Biden. In a well publicized recording of then-president Donald Trump to Georgia Secretary of State Brad Raffensperger, Trump asked Raffensperger to ‘find’ 11,700 votes. [^1] See, e.g., https://int.nyt.com/data/documenttools/highlights-of-trump-s-call-with-the-georgia-secretary-of-state-1/b67c0d9dbde1a697/full.pdf visited 11 July 2024. Subsequently, in early 2021, Trump-affiliated parties gained improper access to all components of the voting system in Coffee County, Georgia and copied and distributed the codebase and data. See, e.g., https://www.cnn.com/2023/08/13/politics/coffee-county-georgia-voting-system-breach-trump/index.html https://apnews.com/article/2022-midterm-elections-technology-lawsuits-donald-trump-voting-6a1324cc6cf45c95ca086a5c81617b15, https://www.washingtonpost.com/investigations/2022/10/28/coffee-county-election-voting-machines/, all accessed 11 July 2024. Georgia hired VotingWorks to orchestrate a ‘risk-limiting audit’ of the 2020 presidential contest, including providing software.
This paper shows that the audit does not support the election results; that the election, recount, and audit disagree; and that all three were unreliable. Among other issues, some memory cards containing votes were not uploaded in the first machine count. Some ballots were included in the first machine tally at least twice. Some ballots were included in the second machine tally at least three times. And some votes manually tabulated in the audit were not included in the reported audit totals. Moreover, the state of Georgia requires all in-person voters to use Dominion Voting Systems (DVS) ballot marking devices (“BMDs”) to mark their ballots. There is extensive evidence that these devices are vulnerable to hacking and misconfiguration [12]; that voters rarely check BMD printout [8, 14, 13]; and that when voters do check, they are unlikely to notice and report printing errors [8, 14]. BMDs are not a trustworthy basis for evidence-based elections [15, 6, 7], even when voted ballots are curated adequately and proper procedures are followed. While there is no evidence of widespread fraud, the mismanagement of the election, reliance on untrustworthy vote records, lack of physical controls on ballots and other voting materials, lack of sanity checks, and poorly executed procedures make it impossible to know who “really” won.
@@ 2 The 2020 audit
Secretary of State Brad Raffensperger claimed, “Georgia’s historic first statewide audit reaffirmed that the state’s new secure paper ballot voting system accurately counted and reported results.” [^2] https://sos.ga.gov/news/historic-first-statewide-audit-paper-ballots-upholds-result-presidential-race, accessed 10 July 2024. Moreover, “[] we did a 100 percent risk-limiting audit with a hand recount which proved the accuracy of the count and also proved that the machines were accurately counting it, and that no votes were flipped.” [^3] https://sos.ga.gov/news/historic-first-statewide-audit-paper-ballots-upholds-result-presidential-race accessed 11 July 2024. VotingWorks Executive Director Ben Adida claimed “Georgia’s first statewide audit successfully confirmed the winner of the chosen contest and should give voters increased confidence in the results.” [^4] Ibid. Per the official report of the audit, “[t]he audit confirmed the original result of the election, namely that Joe Biden won the Presidential Contest in the State of Georgia. The audit [] provides sufficient evidence that the correct winner was reported.” [^5] https://sos.ga.gov/news/historic-first-statewide-audit-paper-ballots-upholds-result-presidential-race accessed 11 July 2024.
@@
Secretary Raffensperger has also used the recount and audit in his defense against a lawsuit that seeks to provide all Georgia voters the option to hand-mark paper ballots in person, rather than being compelled to use BMDs (Curling et al. v. Raffensperger et al., Civil Action No. 1:17-CV-2989-AT, U.S. District Court for the Northern District of Georgia, Atlanta Division). Raffensperger has publicly painted the opposing election security experts in this matter—some of the world’s top cybersecurity experts—as “stop-the-steal” conspiracy theorists, muddying the waters with false claims about the recount and audit and deliberately conflating “there is strong evidence that the election was poorly run and little evidence that the outcome is correct” with “there is strong evidence that the outcome is wrong and that fraud was committed.” Some of the data analyzed below (cast vote records, in particular) were obtained in discovery in Curling v. Raffensperger, but most of the data are a matter of public record and can be downloaded from the Georgia Secretary of State’s website.
@@
The so-called ‘risk-limiting audit’ did not limit the risk of certifying an incorrect electoral outcome for many reasons, starting with its reliance on an untrustworthy record of the votes. The record is untrustworthy because of how it was created (largely BMD printout), curated (a lack of physical accounting for ballots and other materials, lack of pollbook reconciliation, and other elements of a proper canvass), and organized (no “ballot manifest”). The audit could have checked the tabulation of the validly cast ballots it found, but it did not check that properly, as proved by documents on the Secretary of State’s website. [^6] https://sos.ga.gov/news/historic-first-statewide-audit-paper-ballots-upholds-result-presidential-race accessed 11 July 2024.
@@ 2.1 Things the audit did not check
The audit did not check whether BMDs correctly printed voters’ selections. No audit can check that [6]. [^7] As a consequence, Secretary Raffensperger has no basis to assert that no votes were flipped. Expert declarations and testimony in Curling v. Raffensperger establish that the Dominion BMDs can be hacked, misprogrammed, or misconfigured to print votes that differ from voters’ selections as confirmed onscreen or through audio. Only voters are in a position to check—but few do, and those who do check generally check poorly (see citations below). Georgia has no procedures to log, investigate, or report complaints from voters that BMDs altered votes, so it is unknown whether voters observed problems. Logic and accuracy testing cannot establish that BMDs behave correctly in practice [16].
@@
  • The audit did not check whether any scans were duplicated, deleted, replaced or altered.
@@
  • The audit did not aggregate its own results correctly, as explained below.
The exposition below focuses on Fulton County (Atlanta), but there is no reason to believe the problems are confined to Fulton; lapses such as failing to upload memory cards occurred in other counties.
@@ 2.2 The audit report omitted some batch tallies
The audit was conducted using “sort and stack”: teams sorted batches of ballots (including BMD printout) by the presidential vote, then counted the sorted stacks. Batch tallies were manually entered on paper “Audit Board Batch Sheets,’ (ABBSs). Other workers transcribed the ABBSs into VotingWorks audit software “Arlo” to create a database of tallies; totals were calculated from that database. A spreadsheet of results was produced from Arlo. Every ballot validly cast in Fulton County in the 2020 Presidential Election should be reflected in exactly one ABBS, and data from every ABBS should have been (but was not) entered exactly once into the database from which the audit spreadsheet was exported. The transcription of ABBSs was not observable by the public, but the public could in principle compare posted images of the ABBSs to the posted audit spreadsheet, as described below. (Spoiler alert: they do not match.)
Many ABBSs were not completely filled in. The “Batch Type,” signifying the mode of voting (absentee, election day, advance) was often blank, as were many places numbers belonged. The four posted ABBS image files for Fulton County contain a total of 1,927 ABBSs. [^8] Audit subtotals come from the detailed “audit spreadsheet” available at https://sos.ga.gov/admin/uploads/audit-report-November-3-2020-General-Election-2020-11-19.csv (accessed 11 July 2024). Images of the Fulton County, GA, RLA manual tabulation batch sheets were downloaded from https://sos.ga.gov/admin/uploads/Fulton%20RLA%20Batches.zip on 9 January 2022. That file contains five .pdf files, “Fulton Audit Documents 1_redacted.pdf,” through “Fulton Audit Documents 4_redacted.pdf,” which contain images of ABBSs, and “Fulton Audit Documents 5.pdf” which contains images of “Vote Review Panel Tally Sheets.” But the audit spreadsheet contains only 1,916 rows of data for Fulton County. At least eleven ABBSs are entirely missing, not counting possible duplicate entries in the spreadsheet. [^9] However, there is at least one ABBS marked “Dup” (presumably meaning “duplicate”) for instance, page 11 of “Fulton Audit Documents 2_redacted.pdf.” However, as table 1 shows, at least 11 ABBSs are not accounted for in the audit spreadsheet. Thus, there are presumably duplicated entries in the audit spreadsheet. This sort of “sanity check” is simple to perform, but apparently was not performed by the auditors, VotingWorks, Fulton County, or the Secretary of State.
@@ 11 Chastain 114 advance 613 605 24 ? 4 0 3 at 270
Table 1: Examples of audit board batch sheets (ABBSs, tallies of votes in batches of ballots) that were not entered into the audit results spreadsheet.
@@
Vote totals for Donald J. Trump, Joseph R. Biden, and Jo Jorgensen derived by summing ABBS entries in the audit spreadsheet match the vote totals in the summary audit result spreadsheet posted by the Secretary of State at the URL https://sos.ga.gov/admin/uploads/Georgia%202020%20RLA%20Report.xlsx, downloaded on 9 January 2022. The spreadsheet does not list write-ins, undervotes, or overvotes. Both sources show Trump receiving 137,620 votes, Biden receiving 381,179, and Jorgensen receiving 6,494. Thus, the ABBSs that are missing from the audit spreadsheet are also missing from the audit’s reported vote totals.
On the assumption that the ABBSs—the original source of the manual tally data entered into the audit spreadsheet—are correct, the omission of that sample of 11 ABBSs deprived Trump of 1,582 votes, Biden of 2,288, and Jorgensen of 65, not to mention write-ins. This sample alone has a total of over 3,900 votes that the audit tabulated but were not included in the audit’s reported vote totals, compared with a statewide margin of less than 12,000 votes.
@@
There is no way to know whether including those 11 ABBSs makes the audit tabulation a complete count of the votes in Fulton County: many ballots might remain untabulated, because Georgia’s canvass procedures are lax. The proof that at least some of Georgia’s jurisdictions do not keep adequate track of ballots, memory cards, and other election materials is reflected in the fact that thousands of ballots and scans were “discovered” during the audit. [^13] https://www.cbs46.com/news/floyd-county-election-director-fired-after-audit-reveals-2-600-votes-went-uncounted/article_bbd08d90-2aa2-11eb-9e4d-bf96ac56ad54.html, accessed 10 January 2022. https://www.news4jax.com/news/georgia/2020/11/18/4th-georgia-county-finds-uncounted-votes-as-hand-count-deadline-approaches/, accessed 10 January 2022. https://www.mdjonline.com/elections/cobb-elections-finds-350-uncounted-ballots-during-audit/article_0d93e26e-22bd-11eb-8bce-17067aceee33.html, accessed 10 January 2022. https://www.11alive.com/article/news/politics/elections/fayette-county-election-results-ballots-uncovered-during-audit/85-f79dd838-a15c-4407-80b2-9dfbc2466188, accessed 10 January 2022. There is no trustworthy inventory of ballots to check the results against.

### v2

@@
pbstark@berkeley.edu0000-0002-3771-9604
When Audits and Recounts Distract from Election Integrity: The 2020 U.S. Presidential Election in Georgia
To appear in Proceedings of E-Vote-ID 2024, LNI
Philip B. Stark Department of Statistics
University of California
Berkeley, CA 94720-3860
USA
Zusammenfassung
The U.S. state of Georgia was central to efforts to overturn the results of the 2020 Presidential election, including a phone call from then-president Donald Trump to Georgia Secretary of State Brad Raffensperger asking Raffensperger to ‘find’ 11,780 votes. Raffensperger has maintained that a ‘100% full-count risk-limiting audit’ and a machine recount agreed with the initial machine-count results, which proved that the reported election results were accurate and that ‘no votes were flipped.’ While there is no evidence that the reported outcome is wrong, neither is there evidence that it is correct: the two machine counts and the manual ‘audit’ tallies disagree substantially, even about the number of ballots cast. Some ballots in Fulton County, Georgia, were included in the original count at least twice; some were included in the machine recount at least thrice. Audit handcount results for some tally batches were omitted from the reported audit totals: reported audit results do not include all the votes the auditors counted. In short, the two machine counts and the audit were not probative of who won because of poor processes and controls: a lack of secure physical chain of custody, ballot accounting, pollbook reconciliation, and accounting for other election materials such as memory cards. Moreover, most voters used demonstrably untrustworthy ballot-marking devices; as a result, even a perfect handcount or audit would not necessarily reveal who really won. True risk-limiting audits (RLAs) and rigorous recounts can limit the risk that an incorrect electoral outcome will be certified rather than being corrected. But no procedure can limit that risk without a trustworthy record of the vote. And even a properly conducted RLA of some contests in an election does not show that any other contests in that election were decided correctly. The 2020 U.S. Presidential election in Georgia illustrates unrecoverable errors that can render recounts and audits ‘security theater’ that distract from the more serious problems rather than justifying trust.
@@
risk-limiting audit election recount evidence-based elections
1 Introduction: The 2020 U.S. Presidential Election in Georgia
Georgia was one of the ‘swing states’ that determined the outcome of the 2020 U.S. presidential election: its 16 electoral college votes went to Joe Biden. In a well publicized recording of then-president Donald Trump to Georgia Secretary of State Brad Raffensperger, Trump asked Raffensperger to ‘find’ 11,700 votes. [^1] See, e.g., https://int.nyt.com/data/documenttools/highlights-of-trump-s-call-with-the-georgia-secretary-of-state-1/b67c0d9dbde1a697/full.pdf visited 11 July 2024. Subsequently, in early 2021, Trump-affiliated parties gained improper access to all components of the voting system in Coffee County, Georgia and copied and distributed the codebase and data. See, e.g., https://www.cnn.com/2023/08/13/politics/coffee-county-georgia-voting-system-breach-trump/index.html, https://apnews.com/article/2022-midterm-elections-technology-lawsuits-donald-trump-voting-6a1324cc6cf45c95ca086a5c81617b15, https://www.washingtonpost.com/investigations/2022/10/28/coffee-county-election-voting-machines/, all accessed 11 July 2024.
Georgia performed a second machine count and hired VotingWorks to orchestrate a ‘risk-limiting audit’ of the 2020 presidential contest, including providing software. This paper shows that the audit does not support the election results; that the election, recount, and audit disagree; and that all three were unreliable. Among other issues, some memory cards containing votes were not uploaded in the first machine count. Some ballots were included in the first machine tally at least twice. Some ballots were included in the second machine tally at least three times. And some votes manually tabulated in the audit were not included in the reported audit totals. Moreover, the state of Georgia requires all in-person voters to use Dominion Voting Systems (DVS) ballot marking devices (BMDs) to mark their ballots. These devices are vulnerable to hacking and misconfiguration [Ha23]; voters rarely check BMD printout [Be20, KBW20, HI21]; when voters do check, they are unlikely to notice and report printing errors [Be20, KBW20]; and no feasible amount of pre-election testing, logic and accuracy testing, or election-day monitoring can suffice to show that BMDs misbehavior did not alter the outcome [SX22]. BMD printout is thus not a trustworthy basis for evidence-based elections [SW12, ADS20, AS20], even when voted ballots are curated adequately and proper procedures are followed. While there is no evidence of widespread fraud, the mismanagement of the election, reliance on untrustworthy vote records, lack of physical controls on ballots and other voting materials, lack of sanity checks, and poorly executed procedures make it impossible to know who “really” won.
@@ 2 The 2020 audit
Secretary of State Brad Raffensperger claimed, “Georgia’s historic first statewide audit reaffirmed that the state’s new secure paper ballot voting system accurately counted and reported results. …[W]e did a 100 percent risk-limiting audit with a hand recount which proved the accuracy of the count and also proved that the machines were accurately counting it, and that no votes were flipped.” [^2] https://sos.ga.gov/news/historic-first-statewide-audit-paper-ballots-upholds-result-presidential-race, accessed 11 July 2024. VotingWorks Executive Director Ben Adida claimed “Georgia’s first statewide audit successfully confirmed the winner of the chosen contest and should give voters increased confidence in the results.” [^3] Ibid. Per the official report of the audit, “[t]he audit confirmed the original result of the election, namely that Joe Biden won the Presidential Contest in the State of Georgia. The audit [] provides sufficient evidence that the correct winner was reported.” [^4] Ibid.
@@
Secretary Raffensperger has also used the recount and audit in his defense against a lawsuit that seeks to provide all Georgia voters the option to hand-mark paper ballots in person, rather than being compelled to use BMDs (Curling et al. v. Raffensperger et al., Civil Action No. 1:17-CV-2989-AT, U.S. District Court for the Northern District of Georgia, Atlanta Division). Raffensperger has publicly painted the opposing election security experts in this matter—some of the world’s top cybersecurity experts—as “stop-the-steal” conspiracy theorists, muddying the waters with false claims about the recount and audit and deliberately conflating “there is strong evidence that the election was poorly run and little evidence that the outcome is correct” with “there is strong evidence that the outcome is wrong and that fraud was committed.” Some of the data analyzed below (cast vote records, in particular) were obtained in discovery in Curling v. Raffensperger, but most are a matter of public record and can be downloaded from the Georgia Secretary of State’s website, from URLs given below.
@@
The so-called ‘risk-limiting audit’ did not limit the risk of certifying an incorrect electoral outcome for many reasons, starting with its reliance on an untrustworthy record of the votes. The record is untrustworthy because of how it was created (largely BMD printout), curated (a lack of physical accounting for ballots and other materials, lack of pollbook reconciliation, and other elements of a proper canvass), and organized (no “ballot manifest”). The audit could have checked the tabulation of the validly cast ballots it found, but it did not check that properly, as proved by documents on the Secretary of State’s website. [^5] https://sos.ga.gov/news/historic-first-statewide-audit-paper-ballots-upholds-result-presidential-race accessed 11 July 2024. Audit data at the urls https://sos.ga.gov/admin/uploads/Georgia%202020%20RLA%20Report.xlsx, https://sos.ga.gov/admin/uploads/county-summary-data.pdf, and https://sos.ga.gov/admin/uploads/audit-report-November-3-2020-General-Election-2020-11-19.csv, linked from that page, are periodically unavailable, producing the message “Sorry, you have been blocked. You are unable to access sos.ga.gov.” RLA manual tabulation batch sheets were downloaded from https://sos.ga.gov/admin/uploads/Fulton%20RLA%20Batches.zip on 9 January 2022. Precinct-level results for the original machine tally are at https://results.enr.clarityelections.com//GA//105369/271927/reports/detailxls.zip; for the second machine tally, they are at https://results.enr.clarityelections.com//GA//107231/273078/reports/detailxls.zip, both visited 2 September 2024.
@@ 2.1 Things the audit did not check
The audit did not check whether BMDs correctly printed voters’ selections. No audit can check that [ADS20]. (As a consequence, Secretary Raffensperger had no basis to assert that no votes were flipped.) Expert declarations and testimony in Curling v. Raffensperger establish that the Dominion BMDs can be hacked, misprogrammed, or misconfigured to print votes that differ from voters’ selections as confirmed onscreen or through audio. Logic and accuracy testing cannot establish that BMDs behave correctly in practice [SX22]. Only voters are in a position to check—but few do, and those who do check generally check poorly (see citations below). Georgia has no procedures to log, investigate, or report complaints from voters that BMDs altered votes, so it is unknown whether voters observed problems.
@@
  • The audit did not check whether scans were duplicated, deleted, replaced or altered.
@@
  • The audit did not aggregate its own manual tallies correctly, as explained below.
The analysis below focuses on Fulton County (Atlanta), but there is no reason to believe the problems are confined to Fulton; indeed, lapses such as failing to upload memory cards occurred in other counties.
@@ 2.2 The audit report omitted some batch tallies
The audit was conducted using “sort and stack”: teams sorted batches of ballots (including BMD printout) by the presidential vote, then counted the sorted stacks. Batch tallies were manually entered on paper ‘Audit Board Batch Sheets,’ (ABBSs). Other workers transcribed the ABBSs into VotingWorks audit software “Arlo” to create a database of tallies; totals were calculated from that database. A spreadsheet of results was produced from Arlo. Every ballot validly cast in Fulton County in the 2020 Presidential Election should be reflected in exactly one ABBS, and data from every ABBS should have been (but was not) entered exactly once into the database from which the audit spreadsheet was exported. The transcription of ABBSs was not observable by the public, but the public could in principle compare posted images of the ABBSs to the posted audit spreadsheet, as described below. (Spoiler alert: they do not match.)
Many ABBSs were not completely filled in. The “Batch Type,” signifying the mode of voting (absentee, election day, advance) was often blank, as were many places numbers belonged. The four posted ABBS image files for Fulton County contain a total of 1,927 ABBSs. [^6] Audit subtotals come from the detailed “audit spreadsheet” available at https://sos.ga.gov/admin/uploads/audit-report-November-3-2020-General-Election-2020-11-19.csv accessed 11 July 2024. Images of the Fulton County, GA, RLA manual tabulation batch sheets (ABBSs) were downloaded from https://sos.ga.gov/admin/uploads/Fulton%20RLA%20Batches.zip on 9 January 2022. That file contains five .pdf files, “Fulton Audit Documents 1_redacted.pdf,” through “Fulton Audit Documents 4_redacted.pdf,” which contain images of ABBSs, and “Fulton Audit Documents 5.pdf” which contains images of “Vote Review Panel Tally Sheets.” But the audit spreadsheet contains only 1,916 rows of data for Fulton County. At least eleven ABBSs are entirely missing, not counting possible duplicate entries in the spreadsheet. [^7] However, there is at least one ABBS marked “Dup” (presumably meaning “duplicate”) for instance, page 11 of “Fulton Audit Documents 2_redacted.pdf.” However, as table 1 shows, at least 11 ABBSs are not accounted for in the audit spreadsheet. Thus, there are presumably duplicated entries in the audit spreadsheet. This sort of “sanity check” is simple to perform, but apparently was not performed by the auditors, VotingWorks, Fulton County, or the Secretary of State.
@@ 11 Chastain 114 advance 613 605 24 ? 4 0 3 at 270
Tabelle 1: Examples of audit board batch sheets (ABBSs, tallies of votes in batches of ballots) that were not entered into the audit results spreadsheet.
@@
Vote totals for Trump, Biden, and Jorgensen derived by summing ABBS entries in the audit spreadsheet match the vote totals in the summary audit result spreadsheet posted by the Secretary of State at the URL https://sos.ga.gov/admin/uploads/Georgia%202020%20RLA%20Report.xlsx, downloaded on 9 January 2022. The spreadsheet does not list write-ins, undervotes, or overvotes. Both sources show Trump receiving 137,620 votes, Biden receiving 381,179, and Jorgensen receiving 6,494. Thus, the ABBSs that are missing from the audit spreadsheet are also missing from the audit’s reported vote totals.
On the assumption that the ABBSs—the original source of the manual tally data entered into the audit spreadsheet—are correct, the omission of that sample of 11 ABBSs deprived Trump of 1,582 votes, Biden of 2,288, and Jorgensen of 65, not to mention write-ins. This sample alone has a total of over 3,900 votes that the audit tabulated but were not included in Fulton County’s audit vote totals, compared with a statewide margin of less than 12,000 votes.
@@
There is no way to know whether including those 11 ABBSs would make the audit tabulation a complete count in Fulton County: many ballots might remain untabulated, because Georgia’s canvass procedures are lax. The proof some Georgia jurisdictions do not keep adequate track of ballots, memory cards, and other election materials is evidenced by the fact that thousands of ballots and scans were “discovered” during the audit. [^11] https://www.cbs46.com/news/floyd-county-election-director-fired-after-audit-reveals-2-600-votes-went-uncounted/article_bbd08d90-2aa2-11eb-9e4d-bf96ac56ad54.html, accessed 10 January 2022. https://www.news4jax.com/news/georgia/2020/11/18/4th-georgia-county-finds-uncounted-votes-as-hand-count-deadline-approaches/, accessed 10 January 2022. https://www.mdjonline.com/elections/cobb-elections-finds-350-uncounted-ballots-during-audit/article_0d93e26e-22bd-11eb-8bce-17067aceee33.html, accessed 10 January 2022. https://www.11alive.com/article/news/politics/elections/fayette-county-election-results-ballots-uncovered-during-audit/85-f79dd838-a15c-4407-80b2-9dfbc2466188, accessed 10 January 2022. There is no trustworthy inventory of ballots to check the results against.